# On decomposability and subdifferential of the tensor nuclear norm

### v1

@@ 5 Subdifferential of the tensor nuclear norm
and the elements of $\partial f(\mathbf{x})$ are called subgradients; see, e.g., [52, Section 23] and [44, Section 1]. As an immediate consequence of [53, Corollary 8.25], the subdifferential of a norm $\|\bullet\|_{\diamond}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ has a representation
@@
where $\|\bullet\|_{\circ}$ is the dual norm of $\|\bullet\|_{\diamond}$.
@@
see, e.g., [58, Example 2] and [33]. We remark that $\mathbf{U}\mathbf{V}^{\textnormal{T}}$ is invariant over all compact SVDs of $\mathbf{T}$ due to Autonne’s uniqueness [27, Theorem 2.6.5]. The importance of this representation has been well recognized in mathematical optimization and statistics.
In the tensor space, however, only two limited subdifferential inclusions of the nuclear norm are known in the literature, to the best of our knowledge. It was first stated in [61, Lemma 1] that for any third-order tensor $\mathbf{T}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$,
@@
Later, it was stated in [62, Theorem 1] that for any $\mathbf{T}\in\mathbb{R}^{n_{1}\times n_{2}\times\dots\times n_{d}}$,
@@
We remark that although $\mathbb{D}_{2}(\mathbf{T})$ applies to tensors of an arbitrary order, it does not include $\mathbb{D}_{1}(\mathbf{T})$ as a special case. In fact, $\mathbb{D}_{2}(\mathbf{T})$ for $d=3$ is strictly smaller than $\mathbb{D}_{1}(\mathbf{T})$. The two inclusions have been applied in tensor completions [61, 62]. It is also worth mentioning that $\mathbb{Z}(\mathbf{T})$ consists of exactly one element if $\mathbf{T}$ is a matrix, as a consequence of Von Neumann’s trace inequality [51, Theorem 0.1].
@@ 5.2 Subgradients in subspaces
A properly chosen subspace has not only provided better inclusions of the subdifferential of the tensor nuclear norm such as Corollary 5.2 and Theorem 5.3, but also resulted the full decomposability of the tensor nuclear norm such as Theorem 4.4. In this part we look into the details of subgradients of the tensor nuclear norm in various subspaces of interest. In particular, we look into the structure of the subdifferential and estimate the bounds of the spectral norm in relevant subspaces, in order to provide a class of tensors that must be subgradients and a class that cannot be.
@@
    $\operatorname{p}_{\mathbb{T}^{\{1,2\}}(\mathbf{T})}(\mathbf{G})$ can be any matrix as long as $\|\operatorname{p}_{\mathbb{T}^{\{1,2\}}(\mathbf{T})}(\mathbf{G})\|_{\sigma}\leq 1$.
@@
    $\underline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}\right)\geq\frac{1}{2}$ is due to the inclusion $\mathbb{D}_{1}(\mathbf{T})\subseteq\partial\|\mathbf{T}\|_{*}$; see also [61, Lemma 1]. $\underline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}\right)\leq\frac{1}{\sqrt{3}}$ and $\overline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}\right)\geq\frac{2}{\sqrt{3}}$ are due to $\mathbf{Z}+\mathbf{X}(\frac{1}{2})$ and $\mathbf{Z}+\mathbf{X}(-1)$ in Example 5.9, respectively. To upper bound $\overline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}\right)$, we notice that for any $\mathbf{T}$,
@@
    $\underline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[4]}\mathbb{T}^{\mathbb{I}}\right)\geq\frac{1}{3}$ generalizes [61, Lemma 1] from $d=3$ to $d=4$ and its proof is given by Lemma C.2. $\underline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[4]}\mathbb{T}^{\mathbb{I}}\right)\leq\frac{1}{2}$ and $\overline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[4]}\mathbb{T}^{\mathbb{I}}\right)\geq\frac{1+\sqrt{2}}{2}$ are due to the tensors $\mathbf{Z}+\mathbf{X}(\frac{1}{3})$ and $\mathbf{Z}+\mathbf{X}(-\frac{1+\sqrt{2}}{3})$ in Example C.1, respectively. To upper bound $\overline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[4]}\mathbb{T}^{\mathbb{I}}\right)$, we notice that for any $\mathbf{T}$,
@@
    $\underline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[d]}\mathbb{T}^{\mathbb{I}}\right)\geq\frac{2}{d(d-1)}$ is due to the inclusion $\mathbb{D}_{2}(\mathbf{T})\subseteq\partial\|\mathbf{T}\|_{*}$; see also [62, Theorem 1]. Both $\underline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[d]}\mathbb{T}^{\mathbb{I}}\right)\leq\frac{1}{2}$ and $\overline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[d]}\mathbb{T}^{\mathbb{I}}\right)\geq\frac{1+\sqrt{2}}{2}$ are due to (vii) and the monotonicity with respect to $\mathbb{J}$ of $\mathbb{X}^{\mathbb{J}}$ and $d$. Finally, $\overline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[d]}\mathbb{T}^{\mathbb{I}}\right)\leq 2$ is a trivial bound since $\|\mathbf{X}\|_{\sigma}\leq\|\mathbf{X}+\mathbf{Z}\|_{\sigma}+\left\|-\mathbf{Z}\right\|_{\sigma}\leq 2$ for any $\mathbf{Z}\in\mathbb{Z}(\mathbf{T})$ and $\mathbf{X}\in\mathbb{X}^{\mathbb{J}}(\mathbf{T})$ with $\|\mathbf{Z}+\mathbf{X}\|_{\sigma}\leq 1$ in (21).
@@ 5.3 Final remark of the subdifferential
The original version of the subdifferential inclusion in [61, Lemma 1] states that
$\overline{\mathbb{D}}_{1}(\mathbf{T}):=\left\{\mathbf{Z}+\operatorname{p}_{\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}(\mathbf{T})}(\mathbf{X}):\mathbf{Z}\in\mathbb{Z}(\mathbf{T}),\,\mathbf{X}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}},\,\|\mathbf{X}\|_{\sigma}\leq\frac{1}{2}\right\}\subseteq\partial\|\mathbf{T}\|_{*},$
@@
both structures lead to the same set. This is because $\mathbf{X}=\operatorname{p}_{\mathbb{U}^{\mathbb{I}}(\mathbf{T})}(\mathbf{X}+\mathbf{Y})$ and so $\|\mathbf{X}\|_{\sigma}\leq\|\mathbf{X}+\mathbf{Y}\|_{\sigma}\leq 1$ by Lemma 3.2. On the other hand, $\overline{\mathbb{D}}_{1}(\mathbf{T})$ includes $\mathbb{D}_{1}(\mathbf{T})$ as a proper subset since $\|\mathbf{X}\|_{\sigma}\leq\frac{1}{2}$ trivially implies that $\|\mathbf{X}+\mathbf{Y}\|_{\sigma}\leq\frac{1}{2}$ for $\mathbf{Y}=\mathbf{O}$ and the tensor $\mathbf{Z}+\mathbf{X}(\frac{1}{2})$ in Example 5.14 belongs to $\overline{\mathbb{D}}_{1}(\mathbf{T})\setminus\mathbb{D}_{1}(\mathbf{T})$. Even though, both $\overline{\mathbb{D}}_{1}(\mathbf{T})$ and $\mathbb{D}_{1}(\mathbf{T})$ provide the lower bound $\frac{1}{2}$ of $\underline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}\right)$ in Theorem 5.10. In fact, the constant $\frac{1}{2}$ in $\overline{\mathbb{D}}_{1}(\mathbf{T})$ turns out to be tight and this answers in the negative a comment raised in [61, Page 1039] to sharpen the constant.
@@ 6.1 Model and main result
Recall $\mathbb{Z}(\mathbf{L})=\bigl\{\mathbf{Z}\in\mathbb{T}(\mathbf{L}):\langle\mathbf{Z},\mathbf{L}\rangle=\|\mathbf{L}\|_{*},\,\|\mathbf{Z}\|_{\sigma}=1\bigr\}$ defined in (15) for $\mathbf{L}\neq\mathbf{O}$ and the assumption that $2\leq n_{1}\leq n_{2}\leq\dots\leq n_{d}$ without loss of generality. In Assumption 6.1, the first and last requirements of $\mathbf{L}$ are known as the incoherence conditions in the literature [11, 61, 19, 41]. The assumption on $\mathbf{S}$ means that its sparsity patterns follow Bernoulli distributions with parameter $\rho$ but there is no assumption on the magnitudes of the entries of $\mathbf{S}$. We now state the main result in this section.
@@
To the best of our knowledge, Theorem 6.2 is the first result concerning the statistical behavior of (23) for tensors of an arbitrary order. Before proceeding to the proof, let us first compare our result with several existing ones on tensor robust PCA based on nuclear norms. First of all, when $d=2$, the special case of Theorem 6.2 exactly recovers the result on the matrix robust PCA; see [59, Theorem 5.3] and [11, Theorem 1.1].
There are many t-SVD-based methods in the literature that use the t-SVD [31] to define different tensor nuclear norms; see, e.g., [21, 41, 40]. These methods have a good computability and exhibit remarkable performance in numerical experiments but they only work for third-order tensors due to the inherent design mechanism of the t-SVD. This limitation has restricted the generality and versatility of the methods. By contrast, the model (23) is built on top of the vanilla tensor nuclear norm that applies to tensors of an arbitrary order. For third-order tensors, some of the conditions in these t-SVD-based methods seem a bit different from ours. Taking [41, Theorem 4.1] as an example, the balancing parameter $\lambda$ therein is set as $\frac{1}{\sqrt{n_{2}n_{3}}}$ while that in Theorem 6.2 is $\frac{1}{\sqrt{n_{3}}}$, which is consistent with the matrix case $\frac{1}{\sqrt{n_{2}}}$; see [11, Theorem 1.1]. As another example, in [41, Theorem 4.1], the counterpart of $\min_{\mathbf{Z}\in\mathbb{Z}(\mathbf{L})}\|\mathbf{Z}\|_{\infty}$ in its context needs to be bounded by $\sqrt{\frac{u_{\rm{t}}r_{\rm{t}}}{n_{1}n_{2}n_{3}^{2}}}$ where $r_{\rm{t}}$ is the tubal rank of $\mathbf{L}$ [31, Definition 4.4] while ours is $\sqrt{\frac{u_{0}r_{0}}{n_{1}n_{3}\ln{n_{3}}}}$ in Assumption 6.1. The former seems more restrictive if $r_{\rm{t}}$ is regarded analogously to $r_{0}$.
We are also aware of an unpublished work of tensor robust PCA based on the vanilla tensor nuclear norm [19]. We highlight that our result has some advantages. Similar to the t-SVD-based methods, the analysis in [19] also applies to third-order tensors only. Even for third-order tensors, some assumptions in [19] can be stricter than ours. For example, the requirements of $\min_{\mathbf{Z}\in\mathbb{Z}(\mathbf{L})}\|\mathbf{Z}\|_{\infty}$ in [19, Theorem 1] and in Theorem 6.2 are
@@
respectively. While it is difficult to compare the two upper bounds directly, in the case when $r_{k}=r$ and $n_{k}=n$ for $k\in[3]$, they become $O(\sqrt{\frac{r}{n^{3}}})$ and $O(\sqrt{\frac{r}{n^{2}\ln{n}}})$, respectively. It shows that our bound is clearly better in this case. That being said, the rank requirements in [19, Theorem 1] and in Theorem 6.2 are
@@
The rest of this section is devoted to the proof of Theorem 6.2. The overall framework of the proof is similar to that of the matrix robust PCA in [59, Chapter 5] but there are quite a lot of details to be dealt with for tensors.

### v2

@@ 5 Subdifferential of the tensor nuclear norm
and the elements of $\partial f(\mathbf{x})$ are called subgradients; see, e.g., [58, Section 23] and [49, Section 1]. Applying to the tensor nuclear norm, we have for any $\mathbf{T}\in\mathbb{R}^{n_{1}\times n_{2}\times\dots\times n_{d}}$ that
$\partial\|\mathbf{T}\|_{*}=\bigl\{\mathbf{Z}\in\mathbb{R}^{n_{1}\times n_{2}\times\dots\times n_{d}}:\|\mathbf{Y}\|_{*}\geq\|\mathbf{T}\|_{*}+\langle\mathbf{Z},\mathbf{Y}-\mathbf{T}\rangle\,\forall\,\mathbf{Y}\in\mathbb{R}^{n_{1}\times n_{2}\times\dots\times n_{d}}\bigr\}.$
As an immediate consequence of [59, Corollary 8.25], the subdifferential of a norm $\|\bullet\|_{\diamond}:\mathbb{R}^{n}\rightarrow\mathbb{R}$ has a representation
@@
where $\|\bullet\|_{\circ}$ is the dual norm of $\|\bullet\|_{\diamond}$; see (21) for its specialization to the tensor nuclear norm.
@@
see, e.g., [68, Example 2] and [37]. We remark that $\mathbf{U}\mathbf{V}^{\textnormal{T}}$ is invariant over all compact SVDs of $\mathbf{T}$ due to Autonne’s uniqueness [31, Theorem 2.6.5]. The importance of this representation has been well recognized in mathematical optimization and statistics.
In the tensor space, however, only two limited subdifferential inclusions of the nuclear norm are known in the literature, to the best of our knowledge. It was first stated in the discussion following [71, Lemma 1] that for any third-order tensor $\mathbf{T}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}}$,
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\definecolor{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\definecolor{XC@mixtmp}{rgb}{0,0,0}\definecolor{XC@mixtmp}{rgb}{0,0,0}\overline{\mathbb{D}}_{1}(\mathbf{T}):=\left\{\mathbf{Z}+\operatorname{p}_{\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}(\mathbf{T})}(\mathbf{X}):\mathbf{Z}\in\mathbb{Z}(\mathbf{T}),\,\mathbf{X}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}},\,\|\mathbf{X}\|_{\sigma}\leq\frac{1}{2}\right\}\subseteq\partial\|\mathbf{T}\|_{*},$}
@@
Later, it was stated in the discussion following [72, Theorem 1] that for any $\mathbf{T}\in\mathbb{R}^{n_{1}\times n_{2}\times\dots\times n_{d}}$,
@@
In this paper, we work with a slightly different variant of $\overline{\mathbb{D}}_{1}(\mathbf{T})$, i.e.,
which makes the component in $\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}(\mathbf{T})$ explicit. This is mainly to match the same convention in the work [72] for the general order $d$, i.e., an explicit part in $\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[d]}\mathbb{T}^{\mathbb{I}}(\mathbf{T})$ adopted in $\mathbb{D}_{2}(\mathbf{T})$. We believe that this formulation is more natural and transparent; see Section 5.3 for a further discussion. We remark that although $\mathbb{D}_{2}(\mathbf{T})$ applies to tensors of an arbitrary order, it does not include $\mathbb{D}_{1}(\mathbf{T})$ as a special case. In fact, $\mathbb{D}_{2}(\mathbf{T})$ for $d=3$ is strictly smaller than $\mathbb{D}_{1}(\mathbf{T})$. The two inclusions have been applied in tensor completions [71, 72]. It is also worth mentioning that $\mathbb{Z}(\mathbf{T})$ consists of exactly one element if $\mathbf{T}$ is a matrix, as a consequence of Von Neumann’s trace inequality [56, Theorem 0.1].
@@ 5.2 Subgradients in subspaces
A properly chosen subspace has not only provided better inclusions of the subdifferential of the tensor nuclear norm such as Corollary 5.2 and Theorem 5.3, but also resulted in the full decomposability of the tensor nuclear norm such as Theorem 4.4. In this part we look into the details of subgradients of the tensor nuclear norm in various subspaces of interest. In particular, we look into the structure of the subdifferential and estimate the bounds of the spectral norm in relevant subspaces, in order to provide a class of tensors that must be subgradients and a class that cannot be.
@@
    $\operatorname{p}_{\mathbb{T}^{\{1,2\}}(\mathbf{T})}(\mathbf{G})$ can be any matrix in $\mathbb{T}^{\{1,2\}}(\mathbf{T})$ as long as $\|\operatorname{p}_{\mathbb{T}^{\{1,2\}}(\mathbf{T})}(\mathbf{G})\|_{\sigma}\leq 1$.
@@
    $\underline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}\right)\geq\frac{1}{2}$ is due to the inclusion $\mathbb{D}_{1}(\mathbf{T})\subseteq\partial\|\mathbf{T}\|_{*}$; see also [71, Lemma 1]. $\underline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}\right)\leq\frac{1}{\sqrt{3}}$ and $\overline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}\right)\geq\frac{2}{\sqrt{3}}$ are due to $\mathbf{Z}+\mathbf{X}(\frac{1}{2})$ and $\mathbf{Z}+\mathbf{X}(-1)$ in Example 5.9, respectively. To upper bound $\overline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}\right)$, we notice that for any $\mathbf{T}$,
@@
    $\underline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[4]}\mathbb{T}^{\mathbb{I}}\right)\geq\frac{1}{3}$ generalizes [71, Lemma 1] from $d=3$ to $d=4$ and its proof is given by Lemma C.2. $\underline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[4]}\mathbb{T}^{\mathbb{I}}\right)\leq\frac{1}{2}$ and $\overline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[4]}\mathbb{T}^{\mathbb{I}}\right)\geq\frac{1+\sqrt{2}}{2}$ are due to the tensors $\mathbf{Z}+\mathbf{X}(\frac{1}{3})$ and $\mathbf{Z}+\mathbf{X}(-\frac{1+\sqrt{2}}{3})$ in Example C.1, respectively. To upper bound $\overline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[4]}\mathbb{T}^{\mathbb{I}}\right)$, we notice that for any $\mathbf{T}$,
@@
    $\underline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[d]}\mathbb{T}^{\mathbb{I}}\right)\geq\frac{2}{d(d-1)}$ is due to the inclusion $\mathbb{D}_{2}(\mathbf{T})\subseteq\partial\|\mathbf{T}\|_{*}$; see also [72, Theorem 1]. Both $\underline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[d]}\mathbb{T}^{\mathbb{I}}\right)\leq\frac{1}{2}$ and $\overline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[d]}\mathbb{T}^{\mathbb{I}}\right)\geq\frac{1+\sqrt{2}}{2}$ are due to (vii) and the monotonicity with respect to $\mathbb{J}$ of $\mathbb{X}^{\mathbb{J}}$ and $d$. Finally, $\overline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[d]}\mathbb{T}^{\mathbb{I}}\right)\leq 2$ is a trivial bound since $\|\mathbf{X}\|_{\sigma}\leq\|\mathbf{X}+\mathbf{Z}\|_{\sigma}+\left\|-\mathbf{Z}\right\|_{\sigma}\leq 2$ for any $\mathbf{Z}\in\mathbb{Z}(\mathbf{T})$ and $\mathbf{X}\in\mathbb{X}^{\mathbb{J}}(\mathbf{T})$ with $\|\mathbf{Z}+\mathbf{X}\|_{\sigma}\leq 1$ in ($2$).
@@ 5.3 Final remark of the subdifferential
Recall that the original version of the subdifferential inclusion in [71] states that
$\overline{\mathbb{D}}_{1}(\mathbf{T})=\left\{\mathbf{Z}+\operatorname{p}_{\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}(\mathbf{T})}(\mathbf{X}):\mathbf{Z}\in\mathbb{Z}(\mathbf{T}),\,\mathbf{X}\in\mathbb{R}^{n_{1}\times n_{2}\times n_{3}},\,\|\mathbf{X}\|_{\sigma}\leq\frac{1}{2}\right\}\subseteq\partial\|\mathbf{T}\|_{*},$
@@
both structures lead to the same set. This is because $\mathbf{X}=\operatorname{p}_{\mathbb{U}^{\mathbb{I}}(\mathbf{T})}(\mathbf{X}+\mathbf{Y})$ and so $\|\mathbf{X}\|_{\sigma}\leq\|\mathbf{X}+\mathbf{Y}\|_{\sigma}\leq 1$ by Lemma 3.2. On the other hand, $\overline{\mathbb{D}}_{1}(\mathbf{T})$ includes $\mathbb{D}_{1}(\mathbf{T})$ as a proper subset since $\|\mathbf{X}\|_{\sigma}\leq\frac{1}{2}$ trivially implies that $\|\mathbf{X}+\mathbf{Y}\|_{\sigma}\leq\frac{1}{2}$ for $\mathbf{Y}=\mathbf{O}$ and the tensor $\mathbf{Z}+\mathbf{X}(\frac{1}{2})$ in Example 5.14 belongs to $\overline{\mathbb{D}}_{1}(\mathbf{T})\setminus\mathbb{D}_{1}(\mathbf{T})$. Even though, both $\overline{\mathbb{D}}_{1}(\mathbf{T})$ and $\mathbb{D}_{1}(\mathbf{T})$ provide the lower bound $\frac{1}{2}$ of $\underline{\tau}\left(\bigoplus_{|\mathbb{I}|\geq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}\right)$ in Theorem 5.10. In fact, the constant $\frac{1}{2}$ in $\overline{\mathbb{D}}_{1}(\mathbf{T})$ turns out to be tight and this answers in the negative a comment raised in [71, Page 1039] to sharpen the constant.
@@ 6.1 Model and main result
Recall $\mathbb{Z}(\mathbf{L})=\bigl\{\mathbf{Z}\in\mathbb{T}(\mathbf{L}):\langle\mathbf{Z},\mathbf{L}\rangle=\|\mathbf{L}\|_{*},\,\|\mathbf{Z}\|_{\sigma}=1\bigr\}$ defined in (15) for $\mathbf{L}\neq\mathbf{O}$ and the assumption that $2\leq n_{1}\leq n_{2}\leq\dots\leq n_{d}$ without loss of generality. In Assumption 6.1, the first and last requirements of $\mathbf{L}$ are known as the incoherence conditions in the literature [12, 71, 22, 46]. The assumption on $\mathbf{S}$ means that its sparsity patterns follow Bernoulli distributions with parameter $\rho$ but there is no assumption on the magnitudes of the entries of $\mathbf{S}$. We now state the main result in this section, deferring its proof to Section 6.6.
@@
To the best of our knowledge, Theorem 6.2 is the first result concerning the statistical behavior of (24) for tensors of an arbitrary order. Before proceeding to the proof, let us first compare our result with several existing ones on tensor robust PCA based on nuclear norms. First of all, when $d=2$, the special case of Theorem 6.2 exactly recovers the result on the matrix robust PCA; see [69, Theorem 5.3] and [12, Theorem 1.1].
There are many t-SVD-based methods in the literature that use the t-SVD [35] to define different tensor nuclear norms; see, e.g., [24, 46, 45]. These methods have a good computability and exhibit remarkable performance in numerical experiments but they only work for third-order tensors due to the inherent design mechanism of the t-SVD. This limitation has restricted the generality and versatility of the methods. By contrast, the model (24) is built on top of the vanilla tensor nuclear norm that applies to tensors of an arbitrary order. For third-order tensors, some of the conditions in these t-SVD-based methods seem a bit different from ours. Taking [46, Theorem 4.1] as an example, the balancing parameter $\lambda$ therein is set as $\frac{1}{\sqrt{n_{2}n_{3}}}$ while that in Theorem 6.2 is $\frac{1}{\sqrt{n_{3}}}$, which is consistent with the matrix case $\frac{1}{\sqrt{n_{2}}}$; see [12, Theorem 1.1]. As another example, in [46, Theorem 4.1], the counterpart of $\min_{\mathbf{Z}\in\mathbb{Z}(\mathbf{L})}\|\mathbf{Z}\|_{\infty}$ in its context needs to be bounded by $\sqrt{\frac{u_{\rm{t}}r_{\rm{t}}}{n_{1}n_{2}n_{3}^{2}}}$ where $r_{\rm{t}}$ is the tubal rank of $\mathbf{L}$ [35, Definition 4.4] while ours is $\sqrt{\frac{u_{0}r_{0}}{n_{1}n_{3}\ln{n_{3}}}}$ in Assumption 6.1. The former seems more restrictive if $r_{\rm{t}}$ is regarded analogously to $r_{0}$.
We are also aware of an unpublished work of tensor robust PCA based on the vanilla tensor nuclear norm [22]. We highlight that our result has some advantages. Similar to the t-SVD-based methods, the analysis in [22] also applies to third-order tensors only. Even for third-order tensors, some assumptions in [22] can be stricter than ours. For example, the requirements of $\min_{\mathbf{Z}\in\mathbb{Z}(\mathbf{L})}\|\mathbf{Z}\|_{\infty}$ in [22, Theorem 1] and in Theorem 6.2 are
@@
respectively. While it is difficult to compare the two upper bounds directly, in the case when $r_{k}=r$ and $n_{k}=n$ for $k\in[3]$, they become $O(\sqrt{\frac{r}{n^{3}}})$ and $O(\sqrt{\frac{r}{n^{2}\ln{n}}})$, respectively. It shows that our bound is clearly better in this case. That being said, the rank requirements in [22, Theorem 1] and in Theorem 6.2 are
@@
It is worth mentioning that the differences between the recovery bounds required in [22, Theorem 1] and in Theorem 6.2 are driven by different subspaces coupled with different subdifferential inclusions, i.e., (17) and (25). These differences largely govern the quantitative differences in many subsequent tools used in [22] and in this chapter. For example, [22, Lemma 7] and Lemma 6.5 to be introduced soon, both concerning the projections of standard basis tensors, state that
$\max_{\mathbf{i}\in\mathbb{I}^{3}}\left\|\operatorname{p}_{\bigoplus_{|\mathbb{I}|\leq 1,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}(\mathbf{L})}\left(\bigotimes_{k=1}^{3}\mathbf{e}_{i_{k}}\right)\right\|_{2}^{2}\leq\frac{v_{\rm{d}}^{2}r_{\rm{d}}^{2}\sum_{k=1}^{3}n_{k}}{n_{1}n_{2}n_{3}}\mbox{ and }\max_{\mathbf{i}\in\mathbb{I}^{3}}\left\|\operatorname{p}_{\bigoplus_{|\mathbb{I}|\leq 2,\,\mathbb{I}\subseteq[3]}\mathbb{T}^{\mathbb{I}}(\mathbf{L})}\left(\bigotimes_{k=1}^{3}\mathbf{e}_{i_{k}}\right)\right\|_{2}^{2}\leq u_{0}\sum_{k=1}^{3}\frac{r_{k}}{n_{k}},$
respectively. If $r_{\rm{d}}$ is regarded analogously to $r_{0}$ and $n_{k}=n$ for $k\in[3]$, then the above two bounds become $O(\frac{r_{0}^{2}}{n^{2}})$ and $O(\frac{r_{0}}{n})$, respectively. The two bounds scale differently and propagate through subsequent bounds in the analysis, resulting in different recovery bounds. In any case, we believe that the recovery bounds in [22, Theorem 1] and in Theorem 6.2 offer complementary insights, when the order of the tensor is three.
@@
The rest of this section is devoted to the proof of Theorem 6.2. The overall framework of the proof is similar to that of the matrix robust PCA in [69, Chapter 5] but there are quite a lot of details to be dealt with for tensors.